A person walks up at stationary escalator in time t1. If he remains stationary on the escalator,then it can take him up in time t2.How much time would it take him to walk up the moving escalator?

Grade:11

1 Answers

John Sharon Sandeep
6 years ago
This problem can be done using physics speed time relation or mathematical work time (W in unit time= 1/T) relation.Both are same.🔜#Since time taken by mans foot= t1. Therefore steps covered in unit time= 1/t1#And time taken by an escalator step=t2. Therefore steps covered in unit time= 1/t2 🚩thus steps covered in unit time when escalator moves and man too moves= 1/t1 + 1/t2 = (t1+t2)/t1t2.☺thus time taken to complete the distance=1/(t1+t2)/t1t2= t1t2/(t1+t2)